eBook The response to a question has three alternatives: A, B, and C. A sample of 120 responses provides 59 A, 22 B, and 39 C. show the frequency and relative frequency distributions (use 2 decimal for the relative frequency column). Class Frequency Relaive Frequency Total
Answer #1

Relative frequency = Class frequency / Total frequency.

Class Frequency Relative Frequency
A 59 59/120 = 0.49
B 22 22/120 = 0.18
C 39 39/120 = 0.33
TOTAL 120 1.00